# XtalOpt Version 13: Multi-Objective Evolutionary Search for Novel Functional Materials

### v1

@@ 1 Introduction
The previous version of the XtalOpt code [47, 16] included an implementation of the MOES algorithm to search for (meta)stable phases of hard materials. In this scheme, the Vickers hardness, obtained from the AFLOW machine learning (ML) estimated bulk moduli [48], was used in combination with the enthalpy to evaluate the fitness of candidate structures. In addition to finding numerous superhard carbon polymorphs, a superhard-superconducting phase was found [49].
@@ 2.3 Multi-objective search parameters
  • optimization weight: a number between 0.0 and 1.0 used as the weight for the corresponding objective in the fitness calculation. The total weight of all objectives should not exceed 1.0, and (unless otherwise specified) the weight for minimizing the enthalpy is calculated by XtalOpt as: “1.0 - total weight of the objectives”.
@@ 4 Multi-objective runs in the XtalOpt GUI
We now describe how the XtalOpt GUI has been modified to reflect the new options that are available in the MOES run. In the Search Settings tab the AFLOW-ML hardness related entries have been removed (green box in Figure 2(a)), and a new Multiobjective Search tab) is introduced (red outline). In the Multiobjective Search tab (Figure 2(b)), all entries relevant to a MOES run (described in Section 2.3) can be entered in the corresponding fields by (1) choosing the MOES run type from a drop-down menu, (2) specifying the weight associated with a particular objective, entering the (3) name and full path to the user-provided script and (4) output file names, and (5) specifying how XtalOpt should handle a structure that fails a filtration objective (as discussed in Section 6). Clicking the button “Add objective” adds the defined objective to the list of objectives for the run.
@@
  • Desired objectives can be arbitrarily added or removed before the run is started. However, once the search begins, the only MOES-related parameter that can be altered is the one instructing the code how to handle the structures discarded by a “filtration” objective.
@@ 9.3 Termination of the XtalOpt run (CLI mode)
9.4 Run-time log file and debug options (CLI and GUI modes)
-DXTALOPT_DEBUG=ON
Further, if the code is compiled with the
-DMOES_DEBUG=ON

### v2

@@ 1 Introduction
The previous version of the XtalOpt code [47, 16] included an implementation of the MOES algorithm to search for (meta)stable phases of hard materials. In this scheme, the Vickers hardness, obtained from the AFLOW machine learning (ML) estimated shear modulus [48], was used in combination with the enthalpy to evaluate the fitness of candidate structures. In addition to finding numerous superhard carbon polymorphs, a superhard-superconducting phase was found [49].
@@ 2.3 Multi-objective search parameters
  • optimization weight: a number between 0.0 and 1.0 used as the weight for the corresponding objective in the fitness calculation. The total weight of all objectives should not exceed 1.0, and the weight for minimizing the enthalpy is calculated by XtalOpt as: “1.0 - total weight of the objectives”.
@@ 4 Multi-objective runs in the XtalOpt GUI
We now describe how the XtalOpt GUI has been modified to reflect the new options that are available in the MOES run. In the Search Settings tab the AFLOW-ML hardness related entries have been removed (green box in Figure 2(a)), and a new Multiobjective Search tab is introduced (red outline). In the Multiobjective Search tab (Figure 2(b)), all entries relevant to a MOES run (described in Section 2.3) can be entered in the corresponding fields by (1) choosing the MOES run type from a drop-down menu, (2) specifying the weight associated with a particular objective, entering the (3) name and full path to the user-provided script and (4) output file names, and (5) specifying how XtalOpt should handle a structure that fails a filtration objective (as discussed in Section 6). It should be noted that any weight left as zero in the GUI input will result in the corresponding objective being calculated without affecting the optimization. Finally, clicking the button “Add objective” adds the defined objective to the list of objectives for the run.
@@
  • Desired objectives can be arbitrarily added or removed before the run is started. However, once the search begins, the only MOES-related parameter that can be altered is the one instructing the code how to handle the structures discarded by a “filtration” objective (as detailed in Section 6).
@@ 6 Filtering structures: Constrained search
The same can be achieved in the GUI mode by checking the “Handle structures discarded in filtration” box in the Multiobjective Search tab; where the appropriate follow-up action will be taken according to the “If a job fails” entry in the Search Settings tab and similar to the above workflow discussed for the CLI case.
@@ 9.4 New options for the VASP optimizer (CLI and GUI modes)
9.5 Run-time log file and debug options (CLI and GUI modes)
-DXTALOPT_DEBUG=ON
Further, if the code is compiled with the
-DMOES_DEBUG=ON